OpenStudy (angelwings996):

You owe $1,455.69 on a credit card at a 12.8% APR. You pay 15 days after the due date and the late fee is $35.00, making it 45 days of accrued interest. You pay $250.00. How much is your balance after the payment?

OpenStudy (anonymous):

is late fee a monthly fee?

OpenStudy (angelwings996):

That's what I am confused on. I am not sure if it is only a one time fee o you have to do it more than once. I also do not get the 45 days part either

OpenStudy (anonymous):

i not 100% sure but i'd go like this. and it is simplest way i know how 1455.69 +1455.69*.128+35-250= new balance what you owe+ interest + late fee - payment= new balance

OpenStudy (anonymous):

assuming it is all based on a monthly basis

OpenStudy (angelwings996):

What is the first $1455.69 for?

OpenStudy (anonymous):

current balance/ original balance

OpenStudy (angelwings996):

Okay, I get what you are doing, but what is the 45 days for ? Do I need to somehow divide or do something with it? Like I thought maybe 1455.69*12.8%*45/365 or something like tht bcz I remember doing some sort of similar problem like that a while back. Idk just throwing it out there lol

OpenStudy (anonymous):

nah i dont think so. I think 45 days is not enought to constitue 2 months of being late. Assuming it is all monthly based tho

OpenStudy (angelwings996):

Okay so pretty much just do the balance times the interest then add those together...then add the $35 in then subtract the 200?

OpenStudy (anonymous):

sorry i have horrid spelling.

OpenStudy (anonymous):

250

OpenStudy (angelwings996):

250 yeah.. and your fine, I can understand you lol.

OpenStudy (anonymous):

also plz dbl check with a financial guy. I am not the financial math type :P

OpenStudy (angelwings996):

Okay I ill do that. Than you so much for trying to help me!

OpenStudy (anonymous):

dont thank me till you get an A :D

OpenStudy (angelwings996):

lol alright, but thanks for taking the time to try and help me. I appreciate it

jimthompson5910 (jim_thompson5910):

You owe $1,455.69 on a credit card at a 12.8% APR The annual rate is 12.8%, so the monthly rate is 12.8/12 = 1.0666% (roughly) If you start with $1,455.69 as the outstanding balance at the beginning of the month, then 30 days later, you'll have 1455.69 + 1455.69*0.010666*1 = 1471.22 dollars as the outstanding balance (assuming 30 days in a month) However, you don't make the payment until 15 days after the due date, so you've been accruing interest for 30+15 = 45 days. Your outstanding balance will be 1455.69 + 1455.69*0.010666*(45/30) = 1,478.98 dollars. The 45/30 is the time in months Then you add on the late fee of $35 to get 1,478.98+35 = 1,513.98 dollars This is the balance before you make the payment After making the $250 payment your balance drops to 1,513.97 - 250 = 1,263.98 dollars

jimthompson5910 (jim_thompson5910):

The only part that might be tricky is *when* the late fee is applied. It could be applied when the payment is made or when the due date is missed. It's probably when the due date is missed, so you'd accrue interest for 30 days, slap on the late fee, then acrue on 15 more days of interest until you make that $250 payment

OpenStudy (anonymous):

i'd go with jim on this one